# Continual Memorization of Factoids in Large Language Models

### v1

@@ 3.2 Replay Does Not Mitigate Forgetting Fully
Replay-based methods mitigate forgetting by sampling a small portion of data from earlier stages and mixing it with the subsequent dataset during training. Replay from past experience has been a long-established mitigation to prevent forgetting in reinforcement learning research (e.g. Mnih et al., 2013) and more recently continual pretraining, for example for LM Although replay-based methods have proven helpful for continual learning, we hypothesize that they will be less effective for tasks requiring memorization, as the individual instances are largely unrelated (Feldman, 2020; Yang et al., 2023). Table 2 shows that although replay reduces forgetting across the board, the effectiveness is not uniform. Replay has less success to avoid forgetting than non-factoid. We provide full results in §B.2. The replay experiments suggest that manipulating the training dynamics such as exposing the model to different distributions, can affect the model’s ability to recall factoids, even when the replayed factoids are not directly related to the other factoids.
4 REMIX: Random and Generic Data Mixing
4.1 Method
Figure 3: Intuition behind each mixing strategy. Forgetting occurs when $\nabla\mathcal{L}(\theta;D_{A})^{T}\nabla\mathcal{L}(\theta;D_{B})<0$. In stage 1, the model goes from $\theta_{0}$ to $\theta_{A}$ in the parameter space (gray arrow). After stage 2, the model arrives at $\theta_{B}$. The translucent blobs represents low-loss region for each dataset. No Mixing: the opposing angle between the red and blue arrows contributes to forgetting. Mixing at Stage 1: the mixing data protects memorization by shifting the model parameters to reduce the angle between the red and blue arrows while converging to a low loss on $D_{A}$. Mixing at Stage 2: mixing data reduces the interference of $D_{B}$ by lowering the angle between blue and red arrows.
Despite the shortcomings of replay, we make one key observation: when mixing only $10\%$ of the factoids used in stage 1, the post-stage 2 accuracy increases from no mixing at $40.1$ to $83.9\%$ for non-factoid stage 2 (Table 4). This implies the existence of associations that were stored in model weights but could not be retrieved effectively. It is then prudent to ask if these “hidden” associations can be surfaced with a different choice of mixing data.
To answer this question, we propose Random and Generic Data Mixing (REMIX), a forgetting prevention strategy that manipulates the memorization dynamics by mixing extra data into training. The mixing data is sampled from either random word sequences or generic text such as pretraining corpora, which has no overlap with the factoids aiming to memorize in stage 1. Figure 3 illustrates the intuition behind the mixing strategies. For the purpose of developing intuition, we take a drastic simplification to assume the entire optimization is captured by the one-step gradient update. The model $\theta_{0}$ first progresses to $\theta_{A}$ at stage 1 with update: $\theta_{A}=\theta_{0}-\eta\nabla\mathcal{L}(\theta_{0};D_{A})$. Similarly, the model $\theta_{A}$ progresses to $\theta_{B}$ at stage 2: $\theta_{B}=\theta_{A}-\eta\nabla\mathcal{L}(\theta_{A};D_{B})$. In a regular forgetting scenario, the increase in loss after stage 2 is $\mathcal{L}(\theta_{B};D_{A})-\mathcal{L}(\theta_{B};D_{A})$, and can be expanded into $(\theta_{B}-\theta_{A})^{T}\nabla\mathcal{L}(\theta_{A};D_{A})+R=-\eta\nabla\mathcal{L}(\theta;D_{A})^{T}\nabla\mathcal{L}(\theta;D_{B})+R$ where $R$ is the higher-order terms. The first term contributes to forgetting when the two gradients $\nabla\mathcal{L}(\theta;D_{A})$ and $\nabla\mathcal{L}(\theta;D_{B})$ point to opposing directions. With REMIX, $D_{M}$ is mixed into the two stages to prevent forgetting. Specifically, the model progresses in stage 1: $\theta_{A}^{\prime}=\theta_{0}-\eta\nabla\mathcal{L}(\theta_{0};D_{A}\cup D_{M})$ if mixed with $D_{M}$ and progresses to $\theta_{B}^{\prime}=\theta_{A}^{\prime}-\eta\nabla\mathcal{L}(\theta_{0};D_{B}\cup D_{M^{\prime}})$ if mixes with $D_{M^{\prime}}$. REMIX is effective when $\mathcal{L}(\theta_{B}^{\prime};D_{A})<\mathcal{L}(\theta_{B};D_{A})$.
REMIX datasets $D_{M}$.
We explore three data sources for generic data mixing: 1) Knowledge Pile (Fei et al., 2024), 3) Arxiv Pile (Gao et al., 2020), and 4) Fineweb (Penedo et al., 2024). We construct the Random Word Sequence data by collecting a set of uniformly sampled $50$ random word sequences from the NLTK Word Corpus (Bird et al., 2009). We check and ensure no overlap between the factoid data and the mixing data (see details in §B.6). When applying REMIX, we add the mixing data directly to $D_{A}$ in stage 1 and $D_{B}$ in stage 2, therefore the model trains on more data at each stage with mixing. We use Random Word Sequence and Knowledge Pile as the main datasets in the following experiments and later show that other mixing datasets show similar trends. We use $D_{A}:D_{M}=1:2$ and $D_{B}:D_{M}=1:2$ for the main experiments.
4.2 Results
@@
Figure 2 shows the results of factoid tasks with Llama-3-8B. We observe that mixing Random Word Sequences prevents forgetting across the board, improving average accuracy for all $D_{A}$, improving Key-Value Recall ($13.5\%\rightarrow 28.8\%$), PopQA ($47.0\%\rightarrow 53.7\%$), and TriviaQA ($39.8\%\rightarrow 51.0\%$). On the other hand, mixing Knowledge Pile at stage 1 hurts the performance. Mixing at stage 2 shows an opposite trend. We observe drastically better performance with mixing Knowledge Pile, improving the average accuracy for Key-Value Recall ($13.5\%\rightarrow 27.8\%$), PopQA ($47.0\%\rightarrow 86.4\%$), and TriviaQA ($39.8\%\rightarrow 90.6\%$). In contrast, mixing Random Word Sequence at stage 2 exacerbates forgetting. The results align with our prediction that stage 1 mixing relies on data that is unrelated to either $D_{A}$ or $D_{B}$, while stage 2 mixing benefit most when forgetting is severe and the mixing data aligns with $D_{A}$.
@@
Figure 2 shows that the model exhibits consistent results after training on non-factoid data at stage 2. We observe that stage 1 mixing is more beneficial than stage 2 mixing across the board. However, the best mixing data varies for different $D_{A}$. Key-Value Recall benefit most from mixing Random Word Sequence at stage 1 ($38.9\%\rightarrow 77.9\%$), while Knowledge Pile benefit most on PopQA ($58.4\%\rightarrow 82.2\%$) and TriviaQA ($57.9\%\rightarrow 77.8\%$).
@@
Factoid Non-Factoid
LAMA EQ WQ Avg GSM8K MATH EC APPS UC Avg
@@
No Mixing   0.1 15.4 29.6 15.0   4.8   1.5 12.7 13.1 51.9 16.8
Random / K-Pile 47.5 44.1 39.0 43.5 60.1 39.1 52.9 54.8 81.0 57.0
@@
No Mixing 66.9 92.3 89.6 82.9 96.9 96.8 96.9 96.9 96.7 96.8
Random / K-Pile 90.5 92.3 89.0 90.6 91.7 91.6 91.8 91.9 91.3 91.7
@@
No Mixing 71.6 86.4 91.5 83.2   4.8 99.0 95.9 79.9 97.0 75.3
Random / K-Pile 77.0 81.5 83.1 80.5   1.6 91.1 95.3 97.7 90.7 75.3
We report REMIX results for Mistral in Figure 3. For Key-Value Recall, REMIX can successfully prevent forgetting and improve performance after stage 2 training on factoid data ($15.0\%\rightarrow 43.5\%$) and stage 2 training on non-factoid data ($16.8\%\rightarrow 57.0\%$). However, the results are less effective with REMIX since the No Mixing baselines are not affected by forgetting severely to begin with.
@@ 5 Analysis
REMIX learns factoids in earlier layers.
We use Logit Lens (nostalgebraist, 2020) to decode the top $10$ tokens from the representations at each layer using the output embedding. We record the layer index of the first occurrence of the correct token and normalize by the total number of occurrences. This measure indicates how early the correct token first appears. In Figure 4, we compare these mixing stragies 1) No Mixing, 2) Random (stage 1) / K-Pile (stage 2) which successfully prevents forgetting, and 3) K-Pile (stage 1) / None (stage 2) which suffers from forgetting for KVR and TriviaQA. We notice two main differences between the two runs – first, the successful run moves the knowledge to an earlier layer, whereas the unsuccessful one does not change where the factoids are stored. The successful run also diversifies the set of layers that are used, as inferred from the substantial increase in the number of layers that respond to Logit Lens. Both of these changes align with the model protecting the factoids from interference, and support our intuition.
@@
Effect of different mixing data.
Figure 5: Comparison between Knowledge Pile and other generic mixing data sources: ArXiv Pile and FineWeb on Key-Value Recall.
Ablating mixing data length.
Figure 6: Accuracy on Key-Value Recall of varying sequence length with the mixing datasets. Left: Random Word Sequence (mixed at stage 1). Right: Knowledge Pile (mixed at stage 2).
Effect of mixing ratio.
Figure 7: Mixing ratio ablation. x-axis indicates the ratio of the mixing data against the training data. The two left-most plots are both stage 2 mixing (S2) and the right-most two are both stage 1 mixing (S1).
@@
Figure 8: $3$-stage continual memorization setting. $B=*$ refers to the stage 2 task, and $C=*$ refers to the stage 3 task. We use Random mixing at stage 1, K-Pile mixing at stage 2 for WebQA, No Mixing at stage 2 for UltraChat, K-Pile mixing at stage 3 for EntityQA, and No Mixing for MATH at stage 3.
@@ 6 Related Work
Continual learning has been the subject of investigation since early research on connectionist models, which identified catastrophic forgetting as a fundamental challenge (McCloskey & Cohen, 1989; Ratcliff, 1990). Many methods have proposed for mitigating forgetting in continual learning. The simplest approach involves maintaining a memory of examples from previous tasks and replaying them during subsequent training (e.g. Robins, 1995; Chaudhry et al., 2019; Shin et al., 2017). Other methods involve regularization techniques that preserve important weights (e.g. Kirkpatrick et al., 2017; Ke et al., 2023) or reduce the divergence between model predictions (Li & Hoiem, 2017). One group of methods project the gradient for a new task to be orthogonal to the gradients from previous tasks, with the aim of reducing interference between tasks (Lopez-Paz & Ranzato, 2017; Farajtabar et al., 2020). Theoretical analyses (Bennani et al., 2020; Doan et al., 2021) have established that these gradient projection methods mitigate forgetting in the Neural Tangent Kernel regime (Jacot et al., 2018). A number of studies have attempted to characterize the relationship between task similarity and forgetting, empirically and theoretically (Ramasesh et al., 2021; Lee et al., 2021; Evron et al., 2022). See Wang et al. (2023) for a more comprehensive overview. In this paper, we restrict the class of approaches to those that do not change model weights, e.g., via regularization.
Memorization and forgetting in LLMs.
In the context of LLMs, many prior works have investigated the factors that influence memorization during pre-training (Tirumala et al., 2022; Carlini et al., 2023; Mallen et al., 2023; Jagielski et al., 2023). In particular, prior work has observed that instruction tuning can lead to some degradation on general NLP tasks, which has been called an “alignment tax” (Ouyang et al., 2022; Bai et al., 2022). Ouyang et al. (2022) find that this alignment tax can be partly mitigated by mixing pre-training data into the alignment data, and Luo et al. (2023a) find that LLMs forget less when the instruction-tuning data is more diverse.  Kotha et al. (2024) find that fine-tuning LLMs leads to bigger performance degradation on tasks that are more similar to the fine-tuning task (as measured by likelihood under the learned fine-tuning distribution). See Shi et al. (2024) and Wu et al. (2024) for more extensive surveys of continual learning in the context of LLMs.
@@
Our work builds on several recent observations about the effect of fine-tuning an LLM on unfamiliar facts. Kang et al. (2024) find that fine-tuning LLMs on unfamiliar examples (questions that the LLM cannot answer correctly via few-shot prompting) lead the model to “hallucinate” plausible-sounding but incorrect answers to unfamiliar test examples. Similarly, Gekhman et al. (2024) find that unknown examples take longer to learn, and learning unknown examples leads to more hallucination. Ghosal et al. (2024) present a conceptual model for this phenomenon, suggesting that fine-tuning on facts that are weakly represented in the model’s weights can lead the model to pay less attention to entities in the query and instead hallucinate an incorrect response. These studies highlight the difficulty of encoding new facts into a model during fine-tuning. Yang et al. (2024) propose to address this challenge by generating synthetic data for continual pretraining, with the goal of acquiring new knowledge given only a small fine-tuning corpus. This approach can be motivated by mechanistic studies (Allen-Zhu & Li, 2024a; b), which have found that knowledge extraction is possibly only when information appears in diverse forms in the training data (e.g. paraphrases), which leads models to encode information more effectively for later extraction. In contrast, our investigation focuses on the effects of mixing in generic or randomly generated data.
@@
Our work is also related to a line of research aimed at explicitly modifying facts that are encoded in an LLM—for example, to update information about entities to reflect changes in the world (e.g. Zhu et al., 2020; Mitchell et al., 2022; Meng et al., 2022; 2023). Studies have shown that these methods can update individual facts, but do not lead to consistent changes about all of the implications of these updates (Zhong et al., 2023; Cohen et al., 2024). A related line of work has investigated whether specific information can be deliberately removed, or “unlearned,” from neural networks (e.g. Graves et al., 2021; Zhang et al., 2023). Our focus in this paper is on introducing new knowledge while retaining existing knowledge, rather than on modifying or erasing existing knowledge.

### v2

@@ 3.2 Replay Does Not Mitigate Forgetting Fully
Replay-based methods mitigate forgetting by sampling a small portion of data from earlier stages and mixing it with the subsequent dataset during training. Replay from past experience has been a long-established mitigation to prevent forgetting in reinforcement learning research (e.g. Mnih et al., 2013) and more recently continual pretraining for LMs. Although replay-based methods have proven helpful for continual learning, we hypothesize that they will be less effective for tasks requiring memorization, as the individual instances are largely unrelated (Feldman, 2020; Yang et al., 2023). Figure 3 shows that although replay reduces forgetting across the board, the effectiveness is not uniform. Replay has less success to avoid forgetting than non-factoid (full results in §B.3). The experiments suggest that manipulating the training dynamics such as exposing the model to different distributions can affect the model’s ability to recall factoids, even when the replayed factoids are individually independent from other factoids in the same stage.
Figure 3: Replay results averaged across all $D_{B}$.
@@
4 REMIX: Random and Generic Data Mixing
4.1 Method
Despite the shortcomings of replay, we make one key observation: when mixing only $10\%$ of the factoids used in stage 1, the accuracy increases after learning non-factoid data in stage 2 from no mixing at $40.1\%$ to $83.9\%$ (Table 7). This implies the existence of associations that were stored in model weights but could not be retrieved effectively. It is then prudent to ask if these “hidden” associations can be surfaced with a different choice of mixing data. To answer this question, we propose Random and Generic Data Mixing (REMIX), a data mixing strategy that manipulates the memorization dynamics during training to prevent forgetting. The mixing data is sampled from either random word sequences or generic text such as pretraining corpora, which has no overlap with the factoids aiming to memorize in stage 1. Figure 2 illustrates the intuition behind the mixing strategies.
For the purpose of developing intuition, we take the simplification to assume the entire optimization is captured by the one-step gradient update. The model $\mathcal{M}_{0}$ first progresses to $\theta_{A}$ at stage 1 with update: $\theta_{A}=\theta_{0}-\eta\nabla\mathcal{L}(\theta_{0};D_{A})$. Similarly, the model $\theta_{A}$ progresses to $\theta_{B}$ at stage 2: $\theta_{B}=\theta_{A}-\eta\nabla\mathcal{L}(\theta_{A};D_{B})$. In a regular forgetting scenario, the increase in loss after stage 2 is $\mathcal{L}(\theta_{B};D_{A})-\mathcal{L}(\theta_{B};D_{A})$, and can be expanded into $(\theta_{B}-\theta_{A})^{T}\nabla\mathcal{L}(\theta_{A};D_{A})+R=-\eta\nabla\mathcal{L}(\theta;D_{A})^{T}\nabla\mathcal{L}(\theta;D_{B})+R$ where $R$ is the higher-order terms. The first term contributes to forgetting when the two gradients $\nabla\mathcal{L}(\theta;D_{A})$ and $\nabla\mathcal{L}(\theta;D_{B})$ point to opposing directions. With REMIX, $D_{M}$ is mixed into the two stages to prevent forgetting. Specifically, the model progresses in stage 1: $\theta_{A}^{\prime}=\theta_{0}-\eta\nabla\mathcal{L}(\theta_{0};D_{A}\cup D_{M})$ if mixed with $D_{M}$ and progresses to $\theta_{B}^{\prime}=\theta_{A}^{\prime}-\eta\nabla\mathcal{L}(\theta_{0};D_{B}\cup D_{M^{\prime}})$ if mixes with $D_{M^{\prime}}$. REMIX becomes effective when $\mathcal{L}(\theta_{B}^{\prime};D_{A})<\mathcal{L}(\theta_{B};D_{A})$.
REMIX datasets $D_{M}$.
We explore three data sources for generic data mixing: 1) Knowledge Pile (Fei et al., 2024), 3) Arxiv Pile (Gao et al., 2020), and 4) Fineweb (Penedo et al., 2024). We construct the Random Word Sequence data by collecting a set of uniformly sampled $50$ random word sequences from the NLTK Word Corpus (Bird et al., 2009). We check and ensure no overlap between the factoid data and the mixing data (see details in §B.9). When applying REMIX, we add the mixing data directly to $D_{A}$ in stage 1 and $D_{B}$ in stage 2. Therefore, the model trains on more data at each stage with mixing. We use Random Word Sequence and Knowledge Pile as the main datasets in the following experiments and later show that other mixing datasets show similar trends. We use $D_{A}:D_{M}=1:2$ and $D_{B}:D_{M}=1:2$ for the main experiments.
4.2 Results
@@
Table 2 shows the results of factoid tasks with Llama-3-8B. We observe that mixing Random Word Sequences prevents forgetting across the board, improving average accuracy for all $D_{A}$, improving Key-Value Recall ($13.5\%\rightarrow 28.8\%$), PopQA ($47.0\%\rightarrow 53.7\%$), and TriviaQA ($39.8\%\rightarrow 51.0\%$). On the other hand, mixing Knowledge Pile at stage 1 hurts the performance. Mixing at stage 2 shows an opposite trend. We observe drastically better performance with mixing Knowledge Pile, improving the average accuracy for Key-Value Recall ($13.5\%\rightarrow 27.8\%$), PopQA ($47.0\%\rightarrow 86.4\%$), and TriviaQA ($39.8\%\rightarrow 90.6\%$). In contrast, mixing Random Word Sequence at stage 2 exacerbates forgetting. The results align with our prediction that stage 1 mixing relies on data that is unrelated to either $D_{A}$ or $D_{B}$, while stage 2 mixing benefit most when forgetting is severe and the mixing data aligns with $D_{A}$.
@@
Figure 2 shows that the model exhibits consistent results after training on non-factoid data at stage 2. We observe that stage 1 mixing is more beneficial than stage 2 mixing across the board. However, the best mixing data varies for different $D_{A}$. KVR benefit most from mixing Random Word Sequence at stage 1 ($38.9\%\rightarrow 77.9\%$), while Knowledge Pile benefit most on PopQA ($58.4\%\rightarrow 82.2\%$) and TriviaQA ($57.9\%\rightarrow 77.8\%$).
@@
We report REMIX results for Mistral in Table 3 (full results in §B.2). For KVR, REMIX can successfully prevent forgetting and improve performance after stage 2 training on factoid data ($15.0\%\rightarrow 43.5\%$). REMIX’s advantage are less pronounced since the No Mixing baselines are not affected by forgetting too severely.
LAMA EQ WQ Avg
KVR No Mixing   0.1 15.4 29.6 15.0
Random / K-Pile 47.5 44.1 39.0 43.5
PopQA No Mixing 66.9 92.3 89.6 82.9
Random / K-Pile 90.5 92.3 89.0 90.6
TriviaQA No Mixing 71.6 86.4 91.5 83.2
Random / K-Pile 77.0 81.5 83.1 80.5
Table 3: REMIX results for Mistral-7B-v0.3 on Factoid benchmarks. We compare the No Mixing baseline to REMIX that mixes with Random Word Sequence at stage 1 and mixes with Knowledge Pile at stage 2. (EQ = EntityQA, WQ = WebQA.) We provide the complete results in §B.2
Comparison to other baselines.
LAMA EntityQA WebQA Avg
@@
No Mixing   2.1 17.4 33.8 17.8
REMIX (Random / K-Pile) 62.4 69.5 70.2 67.4
Weight Regularization   0.1   4.3 76.7 27.3
Behavior Regularization   0.2 15.6 36.6 17.5
@@
No Mixing   7.7 57.8 72.5 46.0
REMIX (Random / K-Pile) 85.8 90.7 80.5 85.7
Weight Regularization 12.1 67.4 76.7 52.1
Behavior Regularization   7.5 59.3 55.5 40.7
@@
No Mixing   4.3 40.5 68.6 37.8
REMIX (Random / K-Pile) 89.2 89.6 86.5 88.4
Weight Regularization   7.9 58.5 80.3 48.9
Behavior Regularization   6.8 39.0 71.0 38.9
Table 4: Comparison of REMIX to the weight and behavior regularization baselines with the factoid datasets at stage 2.
We compare with two other representative baselines against REMIX Table 4: 1) weight regularization (Kirkpatrick et al., 2017) and 2) behavior regularization (Sun et al., 2020). We use Elastic Weight Consolidation (EWC) for weight regularization and calculate the Fisher score using one backward pass using the current mini-batch for training. For behavior regularization, we add the KL between the training model vs the original reference model to the loss. We observe that the weight regularization baseline and behavior regularization baselines lags behind REMIX by a large margin ($40\%$ on KVR, $30\%$ on PopQA, and $40\%$ on TriviaQA).
@@ 5 Analysis
Robust memorization learns factoids in earlier layers.
We use Logit Lens (nostalgebraist, 2020) to decode the top $10$ tokens from the representations at each layer using the output embedding. We record the layer index of the first occurrence of the correct token, referred to as layer of first occurrence (LoF). LoF is then normalize by the total number of occurrences. This measure indicates how early the correct token first appears. In Figure 4 (left), we compare 1) No Mixing, 2) Random / K-Pile which successfully prevents forgetting, and 3) K-Pile / None which suffers from forgetting for KVR. We notice two main differences between the two runs – first, the successful run moves the knowledge to an earlier layer, whereas the unsuccessful one does not change where the factoids are stored. The successful run also diversifies the set of layers that are used. We aggregate the mean and standard deviation of LoF over 100 examples in Figure 4 (right). The results corroborates with our intuition: the model protecting the factoids from interference when the knowledge is stored earlier (lower mean) and diversified (larger STD) in the layers.
LoF Mean LoF STD Acc
KVR
No Mixing 26.8 15.9 40.6
REMIX (R/K-Pile) 25.9 21.6 80.0
PopQA
No Mixing 23.5 5.1 73.0
REMIX (R/K-Pile) 22.1 5.0 91.0
TriviaQA
No Mixing 24.2 4.7 68.0
REMIX (R/K-Pile) 22.6 5.0 92.0
@@
REMIX enables better knowledge manipulation.
Recent works have shown that manipulating learned knowledge is challenging especially during fine-tuning  (Allen-Zhu & Li, 2024a, b). We design two templates to evaluate: 1) Selective Recall and 2) Recall & Manipulate. For selective recall, the model that has memorized the factoids “X: A, Y: B” needs to answer the question “Here are two keys: X and Y. What is the value of the first key?” with A. For recall & manipulate, the model that has memorized the factoid “XYZ: ABC” needs to answer the question “If the first character in the value of key XYZ is changed to Q, what is the new value of key?” with “QBC”. We show in Table 5 that even though knowledge manipulation remains extremely hard for fine-tuning, REMIX still enables better manipulation of learned knowledge than no mixing, especially on the selective recall template.
Selective Recall Recall & Manipulate
Factoid Non-Factoid Factoid Non-Factoid
No Mixing   0.7   8.6   0.2   1.3
REMIX (R/-)   1.9 29.1   0.8    2.3
REMIX (R/K-Pile) 11.2   8.8    3.4   1.8
Table 5: Knowledge manipulation accuracy on KVR. R = Random Word Sequence. KP = Knowledge Pile. REMIX improves knowledge manipulation over No Mixing.
@@
Other ablation and analysis.
We provide extensive ablations and analysis on the effects of different mixing data, mixing data lengths, ablation of mixing ratio, and REMIX’s impact on downstream tasks in §B.6.
@@ 6 Related Work
Continual learning has been the subject of investigation since early research on connectionist models, which identified catastrophic forgetting as a fundamental challenge (McCloskey & Cohen, 1989; Ratcliff, 1990). Many methods have proposed for mitigating forgetting in continual learning. The simplest approach involves maintaining a memory of examples from previous tasks and replaying them during subsequent training (e.g. Robins, 1995; Chaudhry et al., 2019; Shin et al., 2017). Other methods involve regularization techniques that preserve important weights (e.g. Kirkpatrick et al., 2017; Ke et al., 2023) or reduce the divergence between model predictions (Li & Hoiem, 2017). One group of methods project the gradient for a new task to be orthogonal to the gradients from previous tasks, with the aim of reducing interference between tasks (Lopez-Paz & Ranzato, 2017; Farajtabar et al., 2020). A number of studies have attempted to characterize the relationship between task similarity and forgetting, empirically and theoretically (Ramasesh et al., 2021; Lee et al., 2021; Evron et al., 2022). In this paper, we restrict the class of approaches to those that do not change model weights, e.g., via regularization.
Figure 5: $3$-stage continual memorization setting. $B=*$ refers to the stage 2 task, and $C=*$ refers to the stage 3 task. y-axis refers the accuracy (%) on Key-Value Recall. We use Random mixing at stage 1, K-Pile mixing at stage 2 for WebQA, No Mixing at stage 2 for UltraChat (UC), K-Pile mixing at stage 3 for EntityQA, and No Mixing for MATH at stage 3.
Memorization and forgetting in LMs.
In the context of LMs, many prior works have investigated the factors that influence memorization during pre-training (Tirumala et al., 2022; Carlini et al., 2023; Mallen et al., 2023; Jagielski et al., 2023). In particular, prior work has observed that instruction tuning can lead to some degradation on general NLP tasks, which has been called an “alignment tax” (Ouyang et al., 2022; Bai et al., 2022). Ouyang et al. (2022) find that this alignment tax can be partly mitigated by mixing pre-training data into the alignment data, and Luo et al. (2023a) find that LMs forget less when the instruction-tuning data is more diverse.  Kotha et al. (2024) find that fine-tuning LMs leads to bigger performance degradation on tasks that are more similar to the fine-tuning task (as measured by likelihood under the learned fine-tuning distribution). See Shi et al. (2024) and Wu et al. (2024) for more extensive surveys of continual learning in the context of LMs.
@@
Our work builds on several recent observations about the effect of fine-tuning an LM on unfamiliar facts. Kang et al. (2024) find that fine-tuning LMs on unfamiliar examples (questions that the LM cannot answer correctly via few-shot prompting) lead the model to “hallucinate” plausible-sounding but incorrect answers to unfamiliar test examples. Similarly, Gekhman et al. (2024) find that unknown examples take longer to learn, and learning unknown examples leads to more hallucination. These studies highlight the difficulty of encoding new facts into a model during fine-tuning. Instead of directly learning the facts, Jang et al. (2022) and Seo et al. (2024) study the setting where the facts are embedding in the corpora and need to be learned continually. Yang et al. (2024) propose to address this challenge by generating synthetic data for continual pretraining. This approach can be motivated by mechanistic studies (Allen-Zhu & Li, 2024a, b), which have found that knowledge extraction is possibly only when information appears in diverse forms in the training data (e.g. paraphrases), which leads models to encode information more effectively for later extraction.
@@
Our work is also related to a line of research aimed at explicitly modifying facts that are encoded in an LLM—for example, to update information about entities to reflect changes in the world (e.g. Zhu et al., 2020; Mitchell et al., 2022; Meng et al., 2022, 2023). Studies have shown that these methods can update individual facts, but do not lead to consistent changes about all of the implications of these updates (Zhong et al., 2023; Cohen et al., 2024). A related line of work has investigated whether specific information can be deliberately removed, or “unlearned,” from neural networks (e.g. Graves et al., 2021; Zhang et al., 2023). Our focus in this paper is on introducing new knowledge while retaining existing knowledge, rather than on modifying or erasing existing knowledge.